OpenStudy (kj4uts):

Which of the following are orthogonal to (2, 1)? Check all that apply. A. (1, 2) B. (-2, -3) C. (-3, 6) D. (1, -2) Please explain. Thank you!

OpenStudy (kj4uts):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Two vectors are orthogonal if and only if their dot product is 0 so u dot v = 0 if and only if u and v are orthogonal factors

jimthompson5910 (jim_thompson5910):

u = <a,b> v = <2,1> u dot v = a*2 + b*1 = 2a + b = 0

OpenStudy (kj4uts):

@jim_thompson5910 So then C. (-3,6) and D. (1,-2)

jimthompson5910 (jim_thompson5910):

correct on both